# Dark Matter Substructure or Source Model Systematics? A Case Study of Cluster Lens Abell S1063

### v1

@@ 3.3 Bayesian Nested Sampling
since shapelets involve three non-linear source model parameters. Our model reconstruction is given by
where the lensing operator acts on the source light model to produce the lensed images, which are convolved with the PSF and then pixelized to an appropriate JWST-resolution grid, together denoted by operator, to give realistic image reconstructions. The source model is found through either a shapelets or a Delaunay triangulation reconstruction detailed in Section 3.2. The lensing operator is fully determined by the angular deflections of the Curved Arc Basis model given by Equations (1) and (2). The point spread function operation will depend on the NIRCam filter and determines the resolution of our observed lensed images in each filter. We implement all modeling in this work using lenstronomy (Birrer & Amara 2018, Birrer et al. 2021), a publicly available strong gravitational lensing package for modeling and simulations.
@@ 3.4 Substructure Search
We search for dark matter substructure following the method proposed in Şengül et al. (2023), by exploiting the proximity of potential subhalos to highly magnified regions of the lensed images. If a substructure is present along the line of sight of one of the images, it will cause a perturbation to the gravitational lensing sourced by the cluster, thereby referred to as a perturber. This signal would be picked up by a lens model with a dark matter subhalo on top of the Curved Arc Basis model. Şengül et al. (2023) showed in JWST-quality simulations that the Curved Arc Basis can be used to detect subhalos down to masses of and to measure their concentration, ellipticity, and redshift parameters.
@@ 4 Results
In this section, we present our results of modeling system 4 in Abell S1063 with Curved Arc Basis and performing a dark matter substructure search.
We first fit the lensed images with a smooth Curved Arc Basis model, using shapelet and Delaunay source modeling methods, and find improved agreement across filters with Delaunay, an indication of reduced systematics. Interestingly, our substructure search recovers a subhalo with shapelets but not with Delaunay, illustrating that insufficient source model complexity and inherent source regularization can lead to spurious substructure detections. We highlight the power of multi-band analysis in distinguishing dark matter substructure from source and lens model systematics.
4.1 Smooth Curved Arc Basis
We show the best-fit model reconstructions and residuals of a smooth Curved Arc Basis model obtained with shapelets and Delaunay source modeling methods in Figures 4, 5 and 6, for filters F115W, F150W, and F200W, respectively. The shapelet fits shown use the shapelet order parameters optimized for each filter. Interestingly, the best-fit source models reconstructed using the two methods vary considerably. As perhaps expected from a pixel-based method, Delaunay can reproduce more detail in the complex source light distribution composed of two bright peaks connected with a thin twisted arc. The residuals are lower with Delaunay than with shapelet source modeling, although it is difficult to fairly compare the two fits given the different numbers of degrees of freedom involved in the two source modeling methods.
@@ 4.1.1 Analysis with Large Masks
We first analyze the images in three JWST filters using Large Masks, as shown in Figure 2. The posterior distributions of the Curved Arc Basis parameters obtained with shapelet and Delaunay source modeling methods are shown in Appendix A. Each JWST filter offers an independent realization of the PSF and a different source morphology, but probes the same line-of-sight mass distribution. Therefore, the Curved Arc Basis parameters should agree across different JWST filters, and any tension between lens parameters must arise from systematics in our source or lens modeling. To quantify this tension, we compute for each parameter, where and denote the mean and standard deviation of a parameter measured with the th filter. Our results from both shapelet and Delaunay source modeling methods show a significant average tension between lens parameters from different JWST filters. As discussed in Section 3.1.2, lens modeling systematics might arise from the failure of the local Curved Arc Basis approximation beyond its region of validity. This motivates us to shrink our mask sizes and consider smaller elliptical masks where we expect the Curved Arc Basis approximation to work better.
@@ 4.1.2 Analysis with Small Masks
We repeat the multi-filter analysis with a smooth Curved Arc Basis model using Small Masks (see Figure 2). We show the posterior distributions of the CAB parameters obtained with shapelets and Delaunay source modeling in Appendix B. Reducing mask sizes leads to better agreement between filters with both source modeling methods, although a more significant improvement is seen with Delaunay. This shows that lens modeling systematics played an important role with Large Masks. Hence, multi-band analysis can be critical in determining the region of validity for a local lensing formalism.
Interestingly, Delaunay shows significantly better multi-filter agreement, with average tension between parameters at as compared to with shapelets. Increasing the number of tesselation points to (see Section 3.2.2) in Delaunay source modeling further alleviates the tension to , a improvement from shapelets. However, the average CT scores still point to considerable tension between parameters. Thus, systematic errors dominate the pixel errors, violating the assumption that pixel errors are Gaussian in Equation (7) and thereby biasing the posteriors. This motivates us to boost our JWST errros until the residuals from modeling failure become small compared to the boosted errors, burying systematics under Gaussian noise.
@@ 4.1.3 Artificially Boosted Noise
The tension between filters in our analysis with Small Masks indicates that our modeling systematics cannot fully handle the data quality of the JWST system 4 observations. To prevent systematics from biasing our measured CAB parameters, we boost the JWST error map by a factor of at each pixel. Such boosting of pixel errors was employed by Acebron et al. (2024) to limit the impact of imperfectly modeled areas on the lens model. We show the resulting posterior distributions from shapelet and Delaunay source modeling in Figures 7 and 8, respectively, and compare the means and errors of each parameter in Figure 9. We find that several parameters measured by shapelets are still in tension, with an average between different JWST filters. In contrast, source modeling with Delaunay results in a reduction in the average CT score to , bringing all parameters within a agreement between filters. We note that a two-fold boost in pixel errors leads to a factor of increase in the uncertainties of lens parameters. While error boosting does not solve the root cause of systematics, it allows us to adjust the data quality to a level that our systematics can successfully handle. With boosted errors, Delaunay performs significantly better than shapelets, measuring consistent parameters. To reach the same level of consistency and thereby reliability with shapelets, one would need to further boost pixel errors, reducing the quality of the data and the amount of information we can extract from it.
@@ 4.2 Substructure search
We perform a substructure search using a smooth Curved Arc Basis model and both shapelet and Delaunay source modeling methods. We parametrize the dark matter substructure with an NFW profile and sample for its mass, concentration, and 2D position on the lens plane, assuming a lens redshift for the subhalo (see Section 3.4). In our preliminary substructure search with Large Masks and shapelet source modeling, we used wide uniform priors on all subhalo parameters, listed in Table 3 as Initial Priors. We limited the initial subhalo positional priors to arcseconds, where our coordinates are centered on the brightest image pixels, as proximity to bright regions is important for distinguishing the second-order lensing perturbation caused by subhalos. We searched for substructure in all three images of Abell S1063 system 4 and detected a subhalo in Image #2 with shapelets. In our subsequent multi-filter analysis, we used tighter priors on the subhalo position centered at the preliminary detection location, referred to as Wide and Narrow Priors, listed in Table 3 and illustrated in Figure 10.
@@
We find that source modeling systematics with shapelets can lead to convincing spurious detections of dark matter substructure. In particular, our substructure search with shapelet source modeling results in a detection of a subhalo located at consistent positions in all three filters. And yet, the recovered subhalo concentration differs significantly across the three filters, hinting that the subhalo detection may be spurious. Indeed, these results are not reproduced with Delaunay source modeling, which we find is more robust against spurious substructure detection arising from the inadequate complexity of the source model.
4.2.1 Shapelets
@@
We present the posterior distributions of the subhalo parameters recovered using shapelet source modeling with Large Masks and Small Masks in Figures 11 and 12, respectively. We find that all NFW subhalo parameters are consistent at least between two different JWST filters. However, the concentration parameter measured in F150W consistently hits an upper prior of and is in tension with the posteriors of the F115W and F200W filters.
@@
As discussed in Section 3.2, the complexity of the shapelet source model is controlled by the hyperparameter which should be optimized separately for the smooth and substructure lens models. We show the BIC curves of both models fitted to F115W, F150W, and F200W filters in Appendix C and compare the performance of smooth and substructure models at the optimal of each model. The BIC statistic consistently favors the subhalo model over a smooth CAB model in both F115W and F200 filters. On the other hand, we don’t see a significant difference in the BICs of smooth and substructure models for the F150W filter. It is possible to have a real substructure detection that is not reproduced in every filter if the SNR of those filters is insufficient to detect substructure (Lange et al. 2024). However, our lowest statistical significance for the subhalo is in the filter with an intermediate level of SNR (see Table 1), suggesting that it may be spurious.
@@
We also show the posterior distribution of the subhalo parameters for a range of shapelet order parameters around the optimal in Appendix C. The subhalo parameters are generally consistent across a broad range of shapelet order parameters. This suggests that even the most complex shapelet source models may suffer from inadequate source model complexity or other systematics, leading to a preference for a substructure model. However, in both F115W and F200W filters, the BIC between smooth and perturber models tends to decrease with increasing , another hint that the subhalo detection might be spurious.
@@
We illustrate the detailed statistics of the subhalo detection with shapelets in terms of the BIC and the logarithm of the Bayes factor in Figures 13 and 14, respectively. We repeated our substructure analysis using Small Masks and Large Masks and Wide and Narrow Priors imposed on the position of the subhalo in each of the three filters. With Large Masks and Narrow Priors, we consistently detect a subhalo with decisive statistics of and . Shrinking the positional priors of the subhalo to the region enclosed by Narrow Priors leaves F115W and F200W results unaffected but reduces the statistical significance of the subhalo detection in F150W. This is because the positions recovered with Wide Priors in the intermediate filter disagree with the other two filters, and imposing the subhalo to be near the position detected by the other two filters reduces its statistical significance. Interestingly, switching from Large to Small Masks generally reduces both the and of the subhalo detection in every filter. This suggests that lens modeling systematics arising from the failure of CAB formalism in larger masks also contribute to spurious subhalo detection.
@@ 4.2.2 Delaunay
We show the results of our substructure analysis relying on Delaunay source modeling in Figures 15 and 16. Interestingly, we do not detect a subhalo in any of the three filters, as and consistently favor a smooth model over added substructure in both Large and Small Masks and when using either Narrow or Wide Priors on the subhalo position. This is consistent with our null detection of a subhalo using shapelets with boosted errors. As a result, we see that Delaunay is more robust against spurious subhalo detections. This can be explained by the fact that Delaunay source modeling suffers from lower systematics, as seen by better agreement across filters when using Delaunay instead of shapelet source modeling.
@@ 4.3 Source Regularization
To test whether spurious subhalo detections can result from source regularization, we repeat our substructure search with regularized Delaunay source modeling, which we show in Figures 17 and 18. To enforce some level of regularization in the source, we place linear priors of on the regularization strength parameter. When searching for a subhalo with Narrow Priors on the position, we find that all substructure fits are disfavored by the BIC. Interestingly, however, a wide substructure search reveals a subhalo with in the F200W filter. We conclude that source regularization can indeed lead to spurious subhalo detections.
@@ 4.4 Comparison to Global Model
We present a comparison of our local CAB parameters with the CAB parameters predicted by the global model in Figure 20. The green posteriors show the Curved Arc Basis stretches derived from our analysis of F115W data with Delaunay source modeling. The lilac contours show the and confidence intervals of the CAB parameters derived from the global model. The errors in the 28 sampled global model parameters of the five most massive components were propagated using Monte Carlo simulations of 1,000 random convergence and shear maps. For each map in the sample, we derived the CAB stretches by calculating the average eigenvalues of the lensing Jacobian in cutouts centered at the system 4 images. We note that this procedure underestimates the errors in the CAB parameters derived from the global model, as it does not account for uncertainties in the velocity dispersion measurements of the individual cluster member galaxies.
We find good agreement between the global and local models in all but one of the CAB stretch parameters. Tension between the global and local models is expected, as the former does not incorporate pixel-level information and may be inaccurate at predicting local lensing distortions. At the same time, CAB parameters from local analysis are generally better constrained, even when underestimating the errors in CAB parameters from the global model. These results show that incorporating constraints from CAB in global lens modeling can improve the accuracy of state-of-the-art global cluster lens models. Recent work has made significant progress in incorporating pixel-level information in global cluster models, either by directly modeling bright lensed arcs (Acebron et al. 2024, Broadhurst et al. 2024) or by calculating higher-order moments of multiple images from pixel values without performing additional lens modeling (Pignataro et al. 2021, Bergamini et al. 2021, Diego et al. 2022, Sharon et al. 2022). Summary statistics from Curved Arc Basis analysis can complement these efforts, providing valuable information for global cluster models while maintaining the required computational efficiency to make the problem tractable.

### v2

@@ 3.3 Bayesian Nested Sampling
since Cartesian and elliptical shapelets involve three or five non-linear source model parameters, respectively. Our model reconstruction is given by
where the lensing operator acts on the source light model to produce the lensed images, which are convolved with the PSF and then pixelized to an appropriate JWST-resolution grid, together denoted by operator, to give realistic image reconstructions. The source model is found through either a shapelets or a Delaunay triangulation reconstruction detailed in Section 3.2. The lensing operator is fully determined by the angular deflections of the CAB model given by Equations (1) and (2). The point spread function operation will depend on the NIRCam filter and determines the resolution of our observed lensed images in each filter. We implement all modeling in this work using lenstronomy (Birrer & Amara 2018, Birrer et al. 2021), a publicly available strong gravitational lensing package for modeling and simulations.
@@ 3.4 Substructure Search
We search for dark matter substructure following the method proposed in Şengül et al. (2023), by exploiting the proximity of potential subhalos to highly magnified regions of the lensed images. If a substructure is present along the line of sight of one of the images, it will cause a perturbation to the gravitational lensing sourced by the cluster, thereby referred to as a perturber. This signal would be picked up by a lens model with a dark matter subhalo on top of the CAB model. Şengül et al. (2023) showed in JWST-quality simulations that the CAB can be used to detect subhalos down to masses of and to measure their concentration, ellipticity, and redshift parameters.
@@ 3.4.1 Source Regularization
As described in Section 3.2.2, we perform the regularized source inversion (Eq. 11) at each sampling step using a non-negative least squares optimization. We fit for the regularization parameter together with the lens model parameters in our Bayesian nested sampling framework.
@@ 4 Results
In this section, we present our results of modeling system 4 in Abell S1063 with CAB and performing a dark matter substructure search.
We first fit the lensed images with a smooth CAB model, using shapelet and Delaunay source modeling methods, and find improved agreement across filters with Delaunay, an indication of reduced systematics. Interestingly, our substructure search recovers a subhalo with shapelets but not with Delaunay, illustrating that insufficient source model complexity and inherent source regularization can lead to spurious substructure detections. We highlight the power of multi-band analysis in distinguishing dark matter substructure from source and lens model systematics.
4.1 Smooth CAB
We show the best-fit model reconstructions and residuals of a smooth CAB model obtained with Cartesian shapelets and Delaunay source modeling methods in Figures 4, 5 and 6, for filters F115W, F150W, and F200W, respectively. The Cartesian shapelet fits shown use the shapelet order parameters optimized for each filter. Interestingly, the best-fit source models reconstructed using the two methods vary considerably. As perhaps expected from a pixel-based method, Delaunay can reproduce more detail in the complex source light distribution composed of two bright peaks connected with a thin twisted arc. The residuals are lower with Delaunay than with Cartesian shapelet source modeling, although it is difficult to fairly compare the two fits given the different numbers of degrees of freedom involved in the two source modeling methods.
@@ 4.1.1 Analysis with Large Masks
We first analyze the images in three JWST filters using Large Masks, as shown in Figure 2. The posterior distributions of the CAB parameters obtained with Cartesian shapelet and Delaunay source modeling methods are shown in Appendix A. Each JWST filter offers an independent realization of the PSF and a different source morphology, but probes the same line-of-sight mass distribution. Therefore, the CAB parameters should agree across different JWST filters, and any tension between lens parameters must arise from systematics in our source or lens modeling. To quantify this tension, we compute for each parameter, where and denote the mean and standard deviation of a parameter measured with the th filter. Our results from both Cartesian shapelet and Delaunay source modeling methods show a significant average tension between lens parameters from different JWST filters. As discussed in Section 3.1.2, lens modeling systematics might arise from the failure of the local CAB approximation beyond its region of validity. This motivates us to shrink our mask sizes and consider smaller elliptical masks where we expect the CAB approximation to work better.
@@ 4.1.2 Analysis with Small Masks
We repeat the multi-filter analysis with a smooth CAB model using Small Masks (see Figure 2). We show the posterior distributions of the CAB parameters obtained with Cartesian shapelet and Delaunay source modeling in Appendix B. Reducing mask sizes leads to better agreement between filters with both source modeling methods, although a more significant improvement is seen with Delaunay. This shows that lens modeling systematics played an important role with Large Masks. Hence, multi-band analysis can be critical in determining the region of validity for a local lensing formalism.
Interestingly, Delaunay shows significantly better multi-filter agreement, with average tension between parameters at as compared to with Cartesian shapelets. Increasing the number of tesselation points to (see Section 3.2.2) in Delaunay source modeling further alleviates the tension to , a improvement from Cartesian shapelets. However, the average CT scores still point to considerable tension between parameters. Thus, systematic errors dominate the pixel errors, violating the assumption that pixel errors are Gaussian in Equation (7) and thereby biasing the posteriors. This motivates us to boost our JWST errros until the residuals from modeling failure become small compared to the boosted errors, burying systematics under Gaussian noise.
@@ 4.1.3 Artificially Boosted Noise
The tension between filters in our analysis with Small Masks indicates that our modeling systematics cannot fully handle the data quality of the JWST system 4 observations. To prevent systematics from biasing our measured CAB parameters, we boost the JWST error map by a factor of at each pixel. Such boosting of pixel errors was employed by Acebron et al. (2024) to limit the impact of imperfectly modeled areas on the lens model. We show the resulting posterior distributions from Cartesian shapelet and Delaunay source modeling in Figures 7 and 8, respectively, and compare the means and errors of each parameter in Figure 9. We find that several parameters measured by Cartesian shapelets are still in tension, with an average between different JWST filters. In contrast, source modeling with Delaunay results in a reduction in the average CT score to , bringing all parameters within a agreement between filters. We note that a two-fold boost in pixel errors leads to a factor of increase in the uncertainties of lens parameters. While error boosting does not solve the root cause of systematics, it allows us to adjust the data quality to a level that our systematics can successfully handle. With boosted errors, Delaunay performs significantly better than Cartesian shapelets, measuring consistent parameters. To reach the same level of consistency and thereby reliability with Cartesian shapelets, one would need to further boost pixel errors, reducing the quality of the data and the amount of information we can extract from it.
@@ 4.2 Substructure search
We perform a substructure search using a smooth CAB model and both shapelet and Delaunay source modeling methods. We parametrize the dark matter substructure with an NFW profile and sample for its mass, concentration, and 2D position on the lens plane, assuming a lens redshift for the subhalo (see Section 3.4). In our preliminary substructure search with Large Masks and Cartesian shapelet source modeling, we used wide uniform priors on all subhalo parameters, listed in Table 3 as Initial Priors. We limited the initial subhalo positional priors to arcseconds, where our coordinates are centered on the brightest image pixels, as proximity to bright regions is important for distinguishing the second-order lensing perturbation caused by subhalos. We searched for substructure in all three images of Abell S1063 system 4 and detected a subhalo in Image #2 with Cartesian shapelets. In our subsequent multi-filter analysis, we used tighter priors on the subhalo position centered at the preliminary detection location, referred to as Wide and Narrow Priors, listed in Table 3 and illustrated in Figure 10.
@@
We find that source modeling systematics with shapelets can lead to convincing spurious detections of dark matter substructure. In particular, our substructure search with Cartesian shapelet source modeling results in a detection of a subhalo located at consistent positions in all three filters. And yet, the recovered subhalo concentration differs significantly across the three filters, hinting that the subhalo detection may be spurious. Indeed, these results are not reproduced with Delaunay source modeling, which we find is more robust against spurious substructure detection arising from the inadequate complexity of the source model.
4.2.1 Cartesian shapelets
@@
We present the posterior distributions of the subhalo parameters recovered using Cartesian shapelet source modeling with Large Masks and Small Masks in Figures 11 and 12, respectively. We find that all NFW subhalo parameters are consistent at least between two different JWST filters. However, the concentration parameter measured in F150W consistently hits an upper prior of and is in tension with the posteriors of the F115W and F200W filters.
@@
As discussed in Section 3.2, the complexity of the shapelet source model is controlled by the hyperparameter which should be optimized separately for the smooth and substructure lens models. We show the BIC curves of both models fitted to F115W, F150W, and F200W filters in Appendix C and compare the performance of smooth and substructure models at the optimal of each model. The BIC statistic consistently favors the subhalo model over a smooth CAB model in both F115W and F200 filters. On the other hand, we don’t see a significant difference in the BICs of smooth and substructure models for the F150W filter. It is possible to have a real substructure detection that is not reproduced in every filter if the SNR of those filters is insufficient to detect substructure (Lange et al. 2025). However, our lowest statistical significance for the subhalo is in the filter with an intermediate level of SNR (see Table 1), suggesting that it may be spurious.
@@
We also show the posterior distribution of the subhalo parameters for a range of shapelet order parameters around the optimal in Appendix C. The subhalo parameters are generally consistent across a broad range of shapelet order parameters. This suggests that even the most complex Cartesian shapelet source models may suffer from inadequate source model complexity or other systematics, leading to a preference for a substructure model. However, in both F115W and F200W filters, the BIC between smooth and perturber models tends to decrease with increasing , another hint that the subhalo detection might be spurious.
@@
We illustrate the detailed statistics of the subhalo detection with Cartesian shapelets in terms of the BIC and the logarithm of the Bayes factor in Figures 13 and 14, respectively. We repeated our substructure analysis using Small Masks and Large Masks and Wide and Narrow Priors imposed on the position of the subhalo in each of the three filters. With Large Masks and Narrow Priors, we consistently detect a subhalo with decisive statistics of and . Shrinking the positional priors of the subhalo to the region enclosed by Narrow Priors leaves F115W and F200W results unaffected but reduces the statistical significance of the subhalo detection in F150W. This is because the positions recovered with Wide Priors in the intermediate filter disagree with the other two filters, and imposing the subhalo to be near the position detected by the other two filters reduces its statistical significance. Interestingly, switching from Large to Small Masks generally reduces both the and of the subhalo detection in every filter. This suggests that lens modeling systematics arising from the failure of CAB formalism in larger masks also contribute to spurious subhalo detection.
@@ 4.2.2 Delaunay
We show the results of our substructure analysis relying on Delaunay source modeling in Figures 15 and 16. Interestingly, we do not detect a subhalo in any of the three filters, as and consistently favor a smooth model over added substructure in both Large and Small Masks and when using either Narrow or Wide Priors on the subhalo position. This is consistent with our null detection of a subhalo using Cartesian shapelets with boosted errors. As a result, we see that Delaunay is more robust against spurious subhalo detections. This can be explained by the fact that Delaunay source modeling suffers from lower systematics, as seen by better agreement across filters when using Delaunay instead of shapelet source modeling.
4.2.3 Elliptical Shapelets
We repeat our analysis with shapelets using an elliptical basis for source reconstruction. We show the resulting posterior distributions and subhalo detection statistics in Appendix D. As seen with Cartesian shapelets and Delaunay source modeling, the average CT score in Small Masks () is lower than that of Large Masks (), supporting our conclusion that a growing failure of the local lens model leads to higher systematics in increasing mask sizes. However, switching to an elliptical shapelet basis does not alleviate tension in CAB parameters across filters, as CT scores are comparable between Cartesian and elliptical shapelets. In both Large and Small Masks, a substructure model is favored over a smooth CAB model in only one filter (F200W), with and , respectively. Interestingly, in most filters and masks, the statistical significance of the subhalo detection is weaker as compared to Cartesian shapelets. This suggests that the spurious subhalo is statistically favored partially because of the failure of Cartesian shapelets to capture highly elliptical sources, arising from their implicit regularization that favors more spherical source morphologies. However, Delaunay still outperforms elliptical shapelets in robustness against spurious subhalos and agreement in parameters across filters.
@@ 4.3 Source Regularization
To test whether spurious subhalo detections can result from source regularization, we repeat our substructure search with regularized Delaunay source modeling, which we show in Figures 17 and 18. To enforce some level of regularization in the source, we place linear priors of on the regularization strength parameter. When searching for a subhalo with Narrow Priors on the position, we find that all substructure fits are disfavored by the BIC. Interestingly, however, a wide substructure search reveals a subhalo with in the F200W filter. We conclude that gradient regularization can indeed lead to spurious subhalo detections. This finding is consistent with Nightingale et al. (2018) and Galan et al. (2024), who demonstrate that constant source regularization, where the regularization strength is controlled by a single hyperparameter, can lead to poor reconstructions and biased parameters. They also show that such problems can be resolved with an adaptive regularization scheme (Nightingale et al. 2018).
@@ 4.4 Comparison to Global Model
We present a comparison of our local CAB parameters with the CAB parameters predicted by the global model in Figure 20. Due to a degeneracy to a mass sheet transform, only relative stretch parameters can be constrained by a local lensing analysis with CAB. Hence, we derive the stretch parameters from the global model relative to the first radial stretch, , as constrained by our local analysis. The green posteriors show the CAB stretches derived from our analysis of F115W data with Delaunay source modeling. The lilac contours show the and confidence intervals of the CAB parameters derived from the global model. The errors in the 28 sampled global model parameters of the five most massive components were propagated using Monte Carlo simulations of 1,000 random convergence and shear maps. For each map in the sample, we derived the CAB stretches by calculating the average eigenvalues of the lensing Jacobian in cutouts centered at the system 4 images. We note that this procedure underestimates the errors in the CAB parameters derived from the global model, as it does not account for uncertainties in the velocity dispersion measurements of the individual cluster member galaxies.
We find good agreement between the global and local models in all but one of the CAB stretch parameters. Tension between the global and local models is expected, as the former does not incorporate pixel-level information and may be inaccurate at predicting local lensing distortions. At the same time, CAB parameters from local analysis are generally better constrained, even when underestimating the errors in CAB parameters from the global model. These results show that incorporating constraints from CAB in global lens modeling can improve the accuracy of state-of-the-art global cluster lens models. Recent work has made significant progress in incorporating pixel-level information in global cluster models, either by directly modeling bright lensed arcs (Acebron et al. 2024, Broadhurst et al. 2025) or by calculating higher-order moments of multiple images from pixel values without performing additional lens modeling (Pignataro et al. 2021, Bergamini et al. 2021, Diego et al. 2022, Sharon et al. 2022). Summary statistics from CAB analysis can complement these efforts, providing valuable information for global cluster models while maintaining the required computational efficiency to make the problem tractable.